1.   A perfectly competitive firm produces 100 units of a good and the market price per unit of good is Birr 18. If the given firm is known to earn a normal profit of Birr 6 per unit, how much is the firm’s total cost of production at this level?


1
Expert's answer
2021-09-13T11:21:08-0400

Normal profit is the difference between total revenue and total cost. Total revenue is 18×100=1800, normal profit is 6×100=600, so total cost is 1800-600=1200